OpenStudy (anonymous):

Evaluate the double integral of (y-2x^2)dA over R where R is the region bounded by the square |x|+|y|=1.

OpenStudy (anonymous):

Your square is the picture in the attached file It is bounded by the four lines \[(y=1-x,y=-x-1,y=x+1,y=x-1) \] It should be easy now to set up the integral.

OpenStudy (anonymous):

Your in integral will be \[ \int _{-1}^0\int _{-x-1}^{x+1}\left(y-x^2\right)dydx+\int _0^1\int _{x-1}^{1-x}\left(y-x^2\right)dydx \] compute these two integrals, you will find them equals. Could you have predicted that?

OpenStudy (anonymous):

I have 4 partitions :))

OpenStudy (anonymous):

Only two.

OpenStudy (anonymous):

Do I have to expand all these?

OpenStudy (anonymous):

No, you do integrate first with respect to y, then with respect to x

OpenStudy (anonymous):

I got 2(1/2 -2/3)

OpenStudy (anonymous):

-1/3

OpenStudy (anonymous):

That is right.

OpenStudy (anonymous):

YAAAAAAY! Thanks! You're a genius! :)

OpenStudy (anonymous):

yw